# Learning large softmax mixtures with warm start EM

### v1

@@ B.1 Key lemmas for the proof of Theorem 1
with $c_{0}$ specified in (46). Under the assumptions of Theorem 1, we have, for any ${\bm{\omega}},{\bm{\omega}}^{\prime}\in\mathbb{B}_{d}({\bm{\omega}}^{*},\delta_{0})$ and $k\in[K]$,
@@
$({\bm{\theta}}_{k}-{\bm{\theta}}_{k}^{\prime})^{\top}\left(q_{k}({\bm{\omega}})-q_{k}({\bm{\omega}}^{\prime})\right)\leq-{\mu_{k}\gamma_{k}\over\mu_{k}+\gamma_{k}}\|{\bm{\theta}}_{k}-{\bm{\theta}}_{k}^{\prime}\|_{2}^{2}-{1\over\mu_{k}+\gamma_{k}}\|q_{k}({\bm{\omega}})-q_{k}({\bm{\omega}}^{\prime})\|_{2}^{2}$ (80)
$\mu$-almost surely, as $p\to\infty$.
@@
Under the conditions of Theorem 1, we have, for any ${\bm{\omega}}\in\mathbb{B}_{d}({\bm{\omega}}^{*},\delta_{0})$ ,
@@
$\mu$-almost surely, as $p\to\infty$. Here, for some large constant $C>0$,
$\kappa=CK\overline{\alpha}(1+\overline{\sigma}^{2}\Delta^{2})\exp(-\underline{\sigma}^{2}\Delta^{2}/8)$ (83)
@@
Set $\epsilon_{N}=\sqrt{(L+K)\log(N)/(KN)}$. Under the conditions of Theorem 1, there exists $C>0$ such that
$\mathbb{P}\left\{\left.\sup_{{\bm{\omega}}\in\mathbb{B}_{d}({\bm{\omega}}^{*},\delta_{0})}\max_{k\in[K]}|\widehat{M}_{k}({\bm{\omega}})-M_{k}({\bm{\omega}})|\leq C\epsilon_{N}\right|X_{1},\ldots,X_{p}\right\}\geq 1-\mathcal{O}(N^{-L})$
$\mathbb{P}\left\{\left.\sup_{{\bm{\omega}}\in\mathbb{B}_{d}({\bm{\omega}}^{*},\delta_{0})}\max_{k\in[K]}\|\nabla_{{\bm{\theta}}_{k}}\widehat{Q}({\bm{\omega}}\mid{\bm{\omega}})-\nabla_{{\bm{\theta}}_{k}}Q({\bm{\omega}}\mid{\bm{\omega}})\|_{2}\leq C~{}\overline{\sigma}~{}\epsilon_{N}\right|X_{1},\ldots,X_{p}\right\}\geq 1-\mathcal{O}(N^{-L})$
$\mu$-almost surely, as $p\to\infty$
@@
B.2 Proof of Theorem 1
Proof.
As we pointed out in Section 3.3, at the heart of the analysis of the EM-algorithm lies the simple, well-known principle of Proposition 7. This basic observation shows that it is key to establish (47) with $\phi<1$. The problem at hand is non-standard in that we are dealing with a hybrid between the standard EM for ${\bm{\alpha}}$ in step (42) and a first-order EM for ${\bm{\theta}}_{k}$, $k\in[K]$, in step (3.1). First we study the updates $\widehat{\alpha}_{k}^{(t+1)}$. Let
$M_{k}({\bm{\omega}})$ $=\sum_{j=1}^{p}\pi_{{\bm{\omega}}^{*}}(X_{j})h_{{\bm{\omega}}}({\bm{\theta}}_{k}\mid X_{j})=\alpha_{k}\sum_{j=1}^{p}{\pi_{{\bm{\omega}}^{*}}(X_{j})A_{{\bm{\theta}}_{k}}(X_{j})\over\pi_{\bm{\omega}}(X_{j})}$ (84)
be the population counterpart of $\widehat{M}_{k}({\bm{\omega}})$ in (42). Notice that $M_{k}({\bm{\omega}}^{*})=\alpha_{k}^{*}$ (self-consistency), so we can write
$|\widehat{\alpha}_{k}^{(t+1)}-\alpha_{k}^{*}|$ $=|\widehat{M}_{k}(\widehat{\bm{\omega}}^{(t)})-M_{k}({\bm{\omega}}^{*})|$ (85)
$\leq|\widehat{M}_{k}(\widehat{\bm{\omega}}^{(t)})-M_{k}(\widehat{\bm{\omega}}^{(t)})|+|M_{k}(\widehat{\bm{\omega}}^{(t)})-M_{k}({\bm{\omega}}^{*})|$
$\leq\sup_{d({\bm{\omega}},{\bm{\omega}}^{*})\leq\delta_{0}}|\widehat{M}_{k}({\bm{\omega}})-M_{k}({\bm{\omega}})|+|M_{k}(\widehat{\bm{\omega}}^{(t)})-M_{k}({\bm{\omega}}^{*})|$
The nontrivial Lemmas 4 and 3 imply
$\|\widehat{\bm{\alpha}}^{(t+1)}-{\bm{\alpha}}^{*}\|_{\infty}\leq C\epsilon_{N}+\kappa d(\widehat{\bm{\omega}}^{(t)},{\bm{\omega}}^{*})$
with probability $1-\mathcal{O}(N^{-L})$, with $\epsilon_{N}=\sqrt{(L+K)\log(N)/(KN)}$. For sufficiently large $C_{\Delta}$ in Assumption 7, we can guarantee $\kappa<1$ in Lemma 3.
The analysis for the first-order EM-updates $\widehat{\bm{\theta}}_{k}^{(t+1)}$ is slightly more involved. Recall from (50) that
$Q({\bm{\omega}}\mid{\bm{\omega}}^{\prime})$ $=\sum_{j=1}^{p}\pi_{{\bm{\omega}}^{*}}(X_{j})~{}\sum_{k=1}^{K}h_{{\bm{\omega}}^{\prime}}({\bm{\theta}}_{k}^{\prime}\mid X_{j})\log f_{\bm{\omega}}(X_{j},{\bm{\theta}}_{k})$
$=\sum_{j=1}^{p}\pi_{{\bm{\omega}}^{*}}(X_{j})~{}\sum_{k=1}^{K}{\alpha_{k}^{\prime}A_{{\bm{\theta}}^{\prime}_{k}}(X_{j})\over\pi_{{\bm{\omega}}^{\prime}}(X_{j})}\left[\log(\alpha_{k})+X_{j}^{\top}{\bm{\theta}}_{k}-\log\left(\sum_{\ell=1}^{p}\exp\left({X_{\ell}^{\top}{\bm{\theta}}_{k}}\right)\right)\right].$
is the population level counterpart of $\widehat{Q}({\bm{\omega}}\mid{\bm{\omega}}^{\prime})$ in (40). Its gradient with respect to ${\bm{\theta}}_{k}$ equals
$\nabla_{{\bm{\theta}}_{k}}Q({\bm{\omega}}\mid{\bm{\omega}}^{\prime})=\sum_{j=1}^{p}\pi_{{\bm{\omega}}^{*}}(X_{j})~{}{\alpha_{k}^{\prime}A_{{\bm{\theta}}_{k}^{\prime}}(X_{j})\over\pi_{{\bm{\omega}}^{\prime}}(X_{j})}(X_{j}-\mathbf{X}^{\top}A_{{\bm{\theta}}_{k}}).$
By the definition of the sequence $\widehat{\bm{\theta}}^{(t)}_{k}$, $t\geq 0$, and the triangle inequality, we find
$\leq\|\widehat{\bm{\theta}}^{(t)}_{k}-{\bm{\theta}}_{k}^{*}+\eta_{k}\nabla_{{\bm{\theta}}_{k}}Q(\widehat{\bm{\omega}}^{(t)}\mid{\bm{\omega}}^{*})\|_{2}$
$\qquad+\eta_{k}\|\nabla_{{\bm{\theta}}_{k}}Q(\widehat{\bm{\omega}}^{(t)}\mid\widehat{\bm{\omega}}^{(t)})-\nabla_{{\bm{\theta}}_{k}}Q(\widehat{\bm{\omega}}^{(t)}\mid{\bm{\omega}}^{*})\|_{2}$
By Lemmas 3 and 4, we find
$\|\widehat{\bm{\theta}}_{k}^{(t+1)}-{\bm{\theta}}_{k}^{*}\|_{2}$ $\leq\|\widehat{\bm{\theta}}^{(t)}_{k}-{\bm{\theta}}_{k}^{*}+\eta_{k}q(\widehat{\bm{\omega}}^{(t)})\|_{2}+\eta_{k}\left(\overline{\sigma}\kappa~{}d(\widehat{\bm{\omega}}^{(t)},{\bm{\omega}}^{*})+C\overline{\sigma}\epsilon_{N}\right)$
with probability $1-\mathcal{O}(N^{-L})$. Here, we set $q_{k}({\bm{\omega}}):=\nabla_{{\bm{\theta}}_{k}}Q({\bm{\omega}}\mid{\bm{\omega}}^{*})$. It satisfies $q_{k}({\bm{\omega}}^{*})=0$, and its smoothness and strong-concavity properties are stated in Lemma 2. After we square the first term on the right and work out the squares
Now we make the judicious choice
$\eta_{k}={2\over\mu_{k}+\gamma_{k}}$ (86)
so that the last term cancels and we are left with
$\|\widehat{\bm{\theta}}^{(t)}_{k}-{\bm{\theta}}_{k}^{*}+\eta_{k}q_{k}(\widehat{\bm{\omega}}^{(t)})\|_{2}^{2}$$\leq\left\{1-{4\mu_{k}\gamma_{k}\over(\mu_{k}+\gamma_{k})^{2}}\right\}\|\widehat{\bm{\theta}}^{(t)}_{k}-{\bm{\theta}}_{k}^{*}\|_{2}^{2}$
$={(\mu_{k}-\gamma_{k})^{2}\over(\mu_{k}+\gamma_{k})^{2}}\|\widehat{\bm{\theta}}^{(t)}_{k}-{\bm{\theta}}_{k}^{*}\|_{2}^{2}$
Summarizing, we find with probability $1-\mathcal{O}(N^{-L})$, $\mu$-almost surely, that
where
$\phi$ $=\max_{k}{\mu_{k}-\gamma_{k}\over\mu_{k}+\gamma_{k}}+\kappa\max\left({1\over\underline{\alpha}},\max_{k}{2\overline{\sigma}^{2}\over\mu_{k}+\gamma_{k}}\right)$
$<1-{2(1-2c_{0})\underline{\sigma}^{2}\over(1+c_{0})\overline{\sigma}^{2}+(1-c_{0})\underline{\sigma}^{2}}$
and it follows that $\phi<1$ using $c_{0}<1/2$. We used in the last inequality our 7 to ensure that $\overline{\sigma}^{2}\kappa<c_{0}\underline{\alpha}\,\underline{\sigma}^{2}$. We now have
$d(\widehat{\bm{\omega}}^{(t+1)},{\bm{\omega}}^{*})$ $\leq C{2\over\underline{\alpha}}\epsilon_{N}+\phi~{}d(\widehat{\bm{\omega}}^{(t)},{\bm{\omega}}^{*})$
and since
$\delta_{N}$ $:=C{2\over\underline{\alpha}}\epsilon_{N}\leq(1-2c_{0}){\underline{\sigma}^{2}\over\overline{\sigma}^{2}}\delta_{0}\leq(1-\phi)\delta_{0}$
holds for $N$ large enough, using Assumptions 3, 6 and 7, we conclude our proof. ∎
B.3 Proofs of Lemmas 2, 3 and 4
Throughout the proofs, we use the fact that the event
holds with probability at least $1-p^{-s}$. See, for instance, Lemma 17. This means that this event holds $\mu$-almost surely by the Borel-Cantelli lemma, while condition (46) ensures that
$(\delta_{0}/\underline{\sigma})\|\mathbf{X}\|_{\infty,2}\leq c_{0}$ (88)
holds, $\mu$-almost surely. To prove Lemmas 2, 3 and 4, we first state a few technical lemmas with their proofs given in the end of this section.
@@ B.3 Proofs of Lemmas 2, 3 and 4
The following two lemmas are crucial to the proof of Lemma 3. For any $a\neq k\in[K]$, let $\bar{\bm{\theta}}_{ak}^{*}$ be the midpoint of ${\bm{\theta}}_{a}^{*}$ and ${\bm{\theta}}_{k}^{*}$
@@
Let $\bar{\bm{\theta}}_{ak}^{*}$ be defined in (92), $s\geq 2$ and $\delta>0$. For $p\geq p_{0}(B,s,\overline{\sigma},\delta)$ and $L\leq p^{1-\delta}$, for any $a,k\in[K]$ and any ${\bm{\omega}}\in{\mathbb{B}}_{d}({\bm{\omega}}^{*},\delta_{0})$ with $\delta_{0}$ satisfying (88), we have with probability greater than $1-4p^{1-s}$,
@@
Let ${\bm{\omega}}\in{\mathbb{B}}_{d}({\bm{\omega}}^{*},\delta_{0})$ with $\delta_{0}$ satisfying (88). With probability at least $1-4K^{2}p^{1-s}$, the following holds for any $a,k\in[K]$ with $a\neq k$,
@@
B.3.1 Proof of Lemma 2: strong concavity and smoothness of the oracle criterion $q(\cdot)=Q(\cdot|\omega^{*})$
@@
$q_{k}({\bm{\omega}})=\nabla_{{\bm{\theta}}_{k}}Q({\bm{\omega}}\mid{\bm{\omega}}^{*})$ $=\sum_{j=1}^{p}{\alpha^{*}_{k}A_{{\bm{\theta}}_{k}^{*}}(X_{j})}(X_{j}-\mathbf{X}^{\top}A_{{\bm{\theta}}_{k}})=\alpha_{k}^{*}\mathbf{X}^{\top}(A_{{\bm{\theta}}_{k}^{*}}-A_{{\bm{\theta}}_{k}})$

### v2

@@ B.1 Key lemmas for the proof of Theorem 2
with $c_{0}$ specified in (20). Then for any ${\bm{\omega}},{\bm{\omega}}^{\prime}\in{\mathbb{B}}_{d}({\bm{\omega}}^{*},\delta_{0})$ and $k\in[K]$,
@@
$({\bm{\theta}}_{k}-{\bm{\theta}}_{k}^{\prime})^{\top}\left(q_{k}({\bm{\omega}})-q_{k}({\bm{\omega}}^{\prime})\right)\leq-{\mu_{k}\gamma_{k}\over\mu_{k}+\gamma_{k}}\|{\bm{\theta}}_{k}-{\bm{\theta}}_{k}^{\prime}\|_{2}^{2}-{1\over\mu_{k}+\gamma_{k}}\|q_{k}({\bm{\omega}})-q_{k}({\bm{\omega}}^{\prime})\|_{2}^{2}.$ (67)
@@
Under 1 and (20), we have, for any ${\bm{\omega}}\in{\mathbb{B}}_{d}({\bm{\omega}}^{*},\delta_{0})$ ,
@@
where for some large absolute constant $C>0$,
$\kappa=CK\overline{\alpha}(1+\overline{\sigma}^{2}\Delta^{2})\exp(-\underline{\sigma}^{2}\Delta^{2}/8).$ (70)
@@
Grant 1 and conditions (20) & (25). Set
$\epsilon_{N}=\sqrt{\overline{\alpha}KL\log(N)\over N}.$
There exists some absolute constant $C>0$ such that with probability at least $1-\mathcal{O}(N^{-L})$,
$\sup_{{\bm{\omega}}\in{\mathbb{B}}_{d}({\bm{\omega}}^{*},\delta_{0})}\max_{k\in[K]}\left|\widehat{M}_{k}({\bm{\omega}})-M_{k}({\bm{\omega}})\right|$ $\leq C~\epsilon_{N},$
$\sup_{{\bm{\omega}}\in{\mathbb{B}}_{d}({\bm{\omega}}^{*},\delta_{0})}\max_{k\in[K]}\left\|\nabla_{{\bm{\theta}}_{k}}\widehat{Q}({\bm{\omega}}\mid{\bm{\omega}})-\nabla_{{\bm{\theta}}_{k}}Q({\bm{\omega}}\mid{\bm{\omega}})\right\|_{2}$ $\leq C~\overline{\sigma}~\epsilon_{N}.$
@@
B.2 Proofs of Lemmas 4, 5 and 6
To prove Lemmas 4, 5 and 6, we first state and prove a few technical lemmas.
B.2.1 Technical lemmas used to prove Lemmas 4, 5 and 6
@@
Proof.
This bound, the inequality
[…]
To prove (73), we have
${|\pi_{{\bm{\omega}}}(x_{j})-\pi_{{\bm{\omega}}^{\prime}}(x_{j})|\over\pi_{{\bm{\omega}}}(x_{j})}$
$\leq{1\over\pi_{{\bm{\omega}}}(x_{j})}\sum_{k=1}^{K}\left(|\alpha_{k}-\alpha_{k^{\prime}}|A_{{\bm{\theta}}_{k}}(x_{j})+\alpha_{k}^{\prime}\left|A_{{\bm{\theta}}_{k}}(x_{j})-A_{{\bm{\theta}}_{k}^{\prime}}(x_{j})\right|\right)$
The proof is complete. ∎
The following lemma controls the eigenvalues of $H_{{\bm{\theta}}_{k}}$ defined in (63) for all ${\bm{\theta}}_{k}\in{\mathbb{B}}({\bm{\theta}}_{k}^{*};\delta_{0}/\overline{\sigma})$ with $\delta_{0}$ satisfying (20) and $k\in[K]$.
Lemma 8.
Fix any ${\bm{\theta}}^{*}\in\mathbb{R}^{L}$ and any $\delta_{0}$ satisfying (20). Under (15) and (16), there exists some constant $c=c(c_{0})\in(0,1)$ such that for all $k\in[K]$,
$(1-c)\underline{\sigma}^{2}\leq\lambda_{L}(H_{{\bm{\theta}}_{k}})\leq\lambda_{1}(H_{{\bm{\theta}}_{k}})\leq(1+c)\overline{\sigma}^{2},\qquad\forall~{\bm{\theta}}_{k}\in{\mathbb{B}}({\bm{\theta}}_{k}^{*};\delta_{0}/\overline{\sigma})$ (75)
Proof.
Fix ${\bm{\theta}}^{*}\in\{{\bm{\theta}}_{1}^{*},\ldots,{\bm{\theta}}_{k}^{*}\}$. Write $H_{{\bm{\theta}}^{*}}^{1/2}$ as the matrix square root of $H_{{\bm{\theta}}^{*}}=\mathbf{X}^{\top}\Sigma_{A_{{\bm{\theta}}^{*}}}\mathbf{X}$. Let ${\bm{\theta}}\in{\mathbb{B}}({\bm{\theta}}^{*},\delta_{0}/\overline{\sigma})$ be arbitrary. We first bound from above
$=\sup_{v\in\mathbb{S}^{L-1}}\left(R_{1}(v)+R_{2}(v)\right)$
with
$R_{1}(v)$ $:=\sum_{j=1}^{p}(x_{j}^{\top}H_{{\bm{\theta}}^{*}}^{-1/2}v)^{2}\left(A_{{\bm{\theta}}^{*}}(x_{j})-A_{{\bm{\theta}}}(x_{j})\right),$
$R_{2}(v)$ $:=\left|\sum_{j=1}^{p}x_{j}^{\top}H_{{\bm{\theta}}^{*}}^{-1/2}v\left(A_{{\bm{\theta}}^{*}}(x_{j})+A_{{\bm{\theta}}}(x_{j})\right)\right|~\left|\sum_{j=1}^{p}x_{j}^{\top}H_{{\bm{\theta}}^{*}}^{-1/2}v\left(A_{{\bm{\theta}}^{*}}(x_{j})-A_{{\bm{\theta}}}(x_{j})\right)\right|.$
$\leq{3c_{0}\over\varsigma^{2}}\|H_{{\bm{\theta}}^{*}}^{-1/2}\mathbf{X}^{\top}\mathrm{diag}(A_{{\bm{\theta}}^{*}})\mathbf{X}H_{{\bm{\theta}}^{*}}^{-1/2}\|_{\mathrm{op}}$ by (20)
[…]
Next, we observe that
$R_{2}(v)$ $\leq~2\left|\sum_{j=1}^{p}(x_{j}^{\top}H_{{\bm{\theta}}^{*}}^{-1/2}v)A_{{\bm{\theta}}^{*}}(x_{j})\right|\left|\sum_{j=1}^{p}(x_{j}^{\top}H_{{\bm{\theta}}^{*}}^{-1/2}v)(A_{{\bm{\theta}}^{*}}(x_{j})-A_{{\bm{\theta}}}(x_{j}))\right|$
$\qquad+\left|\sum_{j=1}^{p}(x_{j}^{\top}H_{{\bm{\theta}}^{*}}^{-1/2}v)(A_{{\bm{\theta}}^{*}}(x_{j})-A_{{\bm{\theta}}}(x_{j}))\right|^{2}.$
By repeating the arguments in (76), we find that
$\leq{3c_{0}\over\varsigma^{2}}\Bigl{(}\sum_{j=1}^{p}(x_{j}^{\top}H_{{\bm{\theta}}^{*}}^{-1/2}v)^{2}A_{{\bm{\theta}}^{*}}(x_{j})\Bigr{)}^{1/2}\Bigl{(}\sum_{j=1}^{p}A_{{\bm{\theta}}^{*}}(x_{j})\Bigr{)}^{1/2}$
$\leq{3c_{0}\over\varsigma}\sqrt{\|H_{{\bm{\theta}}^{*}}^{-1/2}\mathbf{X}^{\top}\mathrm{diag}(A_{{\bm{\theta}}^{*}})\mathbf{X}H_{{\bm{\theta}}^{*}}^{-1/2}\|_{\mathrm{op}}}$
$\leq 3c_{0}$ (78)
so that
$\sup_{v\in\mathbb{S}^{L-1}}R_{2}(v)$ $\leq{6c_{0}\over\varsigma}\sup_{v\in\mathbb{S}^{L-1}}\sum_{j=1}^{p}\left|x_{j}^{\top}H_{{\bm{\theta}}^{*}}^{-1/2}v\right|A_{{\bm{\theta}}^{*}}(x_{j})+9c_{0}^{2}\leq 3(2c_{0}+3c_{0}^{2}).$
$\|H_{{\bm{\theta}}^{*}}^{-1/2}(H_{\bm{\theta}}-H_{{\bm{\theta}}^{*}})H_{{\bm{\theta}}^{*}}^{-1/2}\|_{\mathrm{op}}\leq 9c_{0}(1+c_{0}).$
Together with Weyl’s inequality, the eigenvalues of $H_{{\bm{\theta}}^{*}}^{-1/2}H_{\bm{\theta}}H_{{\bm{\theta}}^{*}}^{-1/2}$ satisfy
$\left|1-\lambda_{\ell}(H_{{\bm{\theta}}^{*}}^{-1/2}H_{\bm{\theta}}H_{{\bm{\theta}}^{*}}^{-1/2})\right|\leq 9c_{0}(1+c_{0}),\qquad\forall~1\leq\ell\leq L.$
In particular,
$\left(1-9c_{0}-9c_{0}^{2}\right)\lambda_{L}(H_{{\bm{\theta}}^{*}})\leq\lambda_{L}(H_{\bm{\theta}})\leq\lambda_{1}(H_{\bm{\theta}})\leq\lambda_{1}(H_{{\bm{\theta}}^{*}})\left(1+9c_{0}+9c_{0}^{2}\right)$
which completes the proof. ∎
The following two lemmas are crucial to the proof of Lemma 5. For any $a,k\in[K]$, let $\bar{\bm{\theta}}_{ak}^{*}$ be the midpoint of ${\bm{\theta}}_{a}^{*}$ and ${\bm{\theta}}_{k}^{*}$
@@
For any $a,k\in[K]$ and any ${\bm{\omega}}\in{\mathbb{B}}_{d}({\bm{\omega}}^{*},\delta_{0})$ with $\delta_{0}$ satisfying (20), under conditions (15) and (16), we have
@@
Proof.
${1\over\sum_{\ell=1}^{p}e^{x_{\ell}^{\top}\bar{\bm{\theta}}^{*}}}\sup_{v\in\mathbb{S}^{L-1}}\sum_{j=1}^{p}e^{x_{j}^{\top}\bar{\bm{\theta}}^{*}}v^{\top}\mathbf{X}^{\top}(\mathbf{e}_{j}-A_{{\bm{\theta}}_{k}})(\mathbf{e}_{j}-A_{{\bm{\theta}}_{k}})^{\top}\mathbf{X}v$
$=\sup_{v\in\mathbb{S}^{L-1}}\left\{v^{\top}H_{\bar{\bm{\theta}}^{*}}v+v^{\top}\mathbf{X}^{\top}(A_{{\bm{\theta}}_{k}}-A_{\bar{\bm{\theta}}^{*}})(A_{{\bm{\theta}}_{k}}-A_{\bar{\bm{\theta}}^{*}})^{\top}\mathbf{X}v\right\}$
$\leq\lambda_{1}(H_{\bar{\bm{\theta}}^{*}})+2\|\mathbf{X}^{\top}(A_{{\bm{\theta}}_{k}^{*}}-A_{\bar{\bm{\theta}}^{*}})\|_{2}^{2}+2\|\mathbf{X}^{\top}(A_{{\bm{\theta}}_{k}}-A_{{\bm{\theta}}_{k}^{*}})\|_{2}^{2}$
By repeating the arguments in the proof of (B.2.2), we obtain
$\|\mathbf{X}^{\top}(A_{{\bm{\theta}}_{k}^{*}}-A_{\bar{\bm{\theta}}^{*}})\|_{2}$ $~\leq~\|{\bm{\theta}}_{k}^{*}-\bar{\bm{\theta}}^{*}\|_{2}\sup_{u\in[0,1]}\lambda_{1}(H_{\bar{\bm{\theta}}_{u}^{*}})$
$\|\mathbf{X}^{\top}(A_{{\bm{\theta}}_{k}}-A_{{\bm{\theta}}_{k}^{*}})\|_{2}$ $~\leq~\|{\bm{\theta}}_{k}-{\bm{\theta}}_{k}^{*}\|_{2}\sup_{u\in[0,1]}\lambda_{1}(H_{{\bm{\theta}}_{k,u}})$
where we write $\bar{\bm{\theta}}_{u}^{*}=u{\bm{\theta}}_{a}^{*}+(1-u){\bm{\theta}}_{k}^{*}$ and ${\bm{\theta}}_{k,u}=u{\bm{\theta}}_{k}^{*}+(1-u){\bm{\theta}}_{k}$. The proof is completed by invoking (15) and (75). ∎
Recall that ${\bm{\theta}}_{k,u}=u{\bm{\theta}}_{k}+(1-u){\bm{\theta}}_{k}^{*}$ for any $u\in[0,1]$ and $k\in[K]$.
@@
Let ${\bm{\omega}}\in{\mathbb{B}}_{d}({\bm{\omega}}^{*},\delta_{0})$ with $\delta_{0}$ satisfying (20). Under condition (15), the following holds for any $a,k\in[K]$ with $a\neq k$,
@@
Proof.
[…]
Since $A_{{\bm{\theta}}_{k}^{*}}(x_{j})=e^{x_{j}^{\top}{\bm{\theta}}_{k}^{*}}/N_{{\bm{\theta}}_{k}^{*}}$, we further obtain
$\leq{(\alpha_{a}^{*}\vee\alpha_{k}^{*})~\exp\left(x_{j}^{\top}({\bm{\theta}}_{a}^{*}+{\bm{\theta}}_{k}^{*})/2\right)\over N_{{\bm{\theta}}_{k}^{*}}\exp\left(x_{j}^{\top}({\bm{\theta}}_{a}^{*}-{\bm{\theta}}_{k}^{*})/2\right)+N_{{\bm{\theta}}_{a}^{*}}\exp\left(-x_{j}^{\top}({\bm{\theta}}_{a}^{*}-{\bm{\theta}}_{k}^{*})/2\right)}$
$\leq(\alpha_{a}^{*}\vee\alpha_{k}^{*}){\exp(x_{j}^{\top}\bar{\bm{\theta}}_{ak}^{*})\over 2\sqrt{N_{{\bm{\theta}}_{k}^{*}}N_{{\bm{\theta}}_{a}^{*}}}}$
so that
${\alpha_{a}^{*}\alpha_{k}^{*}\over A_{\bar{\bm{\theta}}_{ak}^{*}}(x_{j})}{A_{{\bm{\theta}}_{a}^{*}}(x_{j})A_{{\bm{\theta}}_{k,u}}(x_{j})\over\pi_{{\bm{\omega}}}(x_{j})}\leq{\alpha_{a}^{*}\vee\alpha_{k}^{*}\over 2}{N_{\bar{\bm{\theta}}_{ak}^{*}}\over\sqrt{N_{{\bm{\theta}}_{k}^{*}}N_{{\bm{\theta}}_{a}^{*}}}}.$ (80)
It remains to bound from above
$\log N_{\bar{\bm{\theta}}_{ak}^{*}}-{1\over 2}\left(\log N_{{\bm{\theta}}_{k}^{*}}+\log N_{{\bm{\theta}}_{a}^{*}}\right).$
By letting $g({\bm{\theta}})=\log N_{{\bm{\theta}}}=\log(\sum_{j=1}^{p}e^{x_{j}^{\top}{\bm{\theta}}})$ for any ${\bm{\theta}}\in\mathbb{R}^{L}$, if there exists some $\nu>0$ such that $g({\bm{\theta}})$ is strongly $\nu$-convex over $u{\bm{\theta}}_{a}^{*}+(1-u){\bm{\theta}}_{k}^{*}$ for all $u\in[0,1]$, then
$\log N_{\bar{\bm{\theta}}_{ak}^{*}}-{1\over 2}\left(\log N_{{\bm{\theta}}_{k}^{*}}+\log N_{{\bm{\theta}}_{a}^{*}}\right)=g(\bar{\bm{\theta}}_{ak}^{*})-{1\over 2}\left[g({\bm{\theta}}_{a}^{*})+g({\bm{\theta}}_{k}^{*})\right]\leq-{\nu\over 8}\|{\bm{\theta}}_{a}^{*}-{\bm{\theta}}_{k}^{*}\|_{2}^{2}$
which yields the desired result. To verify the strongly $\nu$-convexity of $g({\bm{\theta}})$, taking the derivative with respect to ${\bm{\theta}}$ twice and interchanging the expectation with derivatives give
$\nabla^{2}g({\bm{\theta}})={\sum_{j=1}^{p}x_{j}x_{j}^{\top}e^{x_{j}^{\top}{\bm{\theta}}}\over\sum_{\ell=1}^{p}e^{x_{\ell}^{\top}{\bm{\theta}}}}-{(\sum_{j=1}^{p}x_{j}e^{x_{j}^{\top}{\bm{\theta}}})(\sum_{j=1}^{p}x_{j}e^{x_{j}^{\top}{\bm{\theta}}})^{\top}\over(\sum_{\ell=1}^{p}e^{x_{\ell}^{\top}{\bm{\theta}}})^{2}}=H_{{\bm{\theta}}}.$
By condition (15), we know that $\lambda_{L}(H_{{\bm{\theta}}})\geq\underline{\sigma}^{2}$ for all ${\bm{\theta}}=u{\bm{\theta}}_{a}^{*}+(1-u){\bm{\theta}}_{k}^{*}$. This implies the strong $\underline{\sigma}^{2}$-convexity hence completes the proof. ∎
B.2.2 Proof of Lemma 4: strong concavity and smoothness of the gradient function $q(\cdot)=\nabla_{\theta_{k}}Q(\cdot\mid\omega^{*})$
@@
$q_{k}({\bm{\omega}})=\nabla_{{\bm{\theta}}_{k}}Q({\bm{\omega}}\mid{\bm{\omega}}^{*})$ $=\sum_{j=1}^{p}{\alpha^{*}_{k}A_{{\bm{\theta}}_{k}^{*}}(x_{j})}(x_{j}-\mathbf{X}^{\top}A_{{\bm{\theta}}_{k}})=\alpha_{k}^{*}~\mathbf{X}^{\top}(A_{{\bm{\theta}}_{k}^{*}}-A_{{\bm{\theta}}_{k}})$